The length of a rectangular garden is 3 meters more than seven times its width. Find the length and width of the garden if the perimeter of the garden is 70 meters.
Answer by checkley77(12569)

L=7W+3
P=2L+2W
70=2(7W+3)+2W
70=14W+6+2W
70=16W+6
16W=70-6
16W=64
W=64/16
W=4 M. ANS. FOR THE WIDTH.
L=7*4+3
L=28+3
L=31 M. ANS FOR THE LENGTH.
PROOF:
70=2*31+2*4
70=62+8
70=70